OpenStudy (anonymous):

solve this equation for x and round to the nearest 100th 1=In(x+4)

OpenStudy (welshfella):

x + 4 = e^1 = e

OpenStudy (welshfella):

solve for x

OpenStudy (anonymous):

-4?

OpenStudy (welshfella):

no x + 4 = e not 0

OpenStudy (anonymous):

-1.28

OpenStudy (welshfella):

ln means log to the base e you can find the value of e on your calculator

OpenStudy (anonymous):

i used it

OpenStudy (welshfella):

looks good

OpenStudy (welshfella):

recall what logs are